# Renaissance: Investigating the Pretraining of Vision-Language Encoders

### v1

@@ Abstract
In the past several years there has been an explosion of available models for vision-language tasks. Unfortunately, the literature still leaves open a number of questions related to best practices in designing and training such models. In this paper we seek to answer several questions related to the pretraining of vision-language encoders through meta-analysis. In our first set of experiments, we show that we can save significant compute at no cost to downstream performance, by freezing large parts of vision-language models during pretraining. In our second set of experiments we examine the effect of basing a VL transformer on a vision model versus a text model. Additionally, we introduce a VL modeling platform called Renaissance that we use to conduct all of the experiments. This program offers a great deal of flexibility in creating, training and evaluating transformer encoders for VL modeling. The source code for Renaissance can be found at https://github.com/bsu-slim/renaissance.
1 Introduction
@@
In the span of a few years, dozens of vision-language (VL) transformers have appeared in the literature with a bewildering array of architectures and training methods (see Fields and Kennington (2023) for a review). VL tasks, such as NLVR2 Suhr et al. (2018) where the model is tasked with answering questions about images (see Figure 4 for an example) and image captioning require models to somehow represent and fuse both text and image information. Unfortunately, knowledge of best practices for training and implementing these models has lagged far behind the model development process. This stands in contrast to the NLP domain, where studies such as Rogers et al. (2021) and Kaplan et al. (2020) have thoroughly investigated the inner workings and best training practices for NLP transformers. To date, there have been only a handful of studies analyzing VL-transformers, such as Bugliarello et al. (2021), and the collected literature still fails to address some very basic questions concerning VL modeling with transformers.
@@
In this paper we begin to address this gap by providing a systematic analysis geared toward shedding light on some basic aspects of training transformers for vision-language modeling. In particular, we focus on the pretraining and fine-tuning of transformer-encoder architectures. Transformer encoders are best suited toward discriminative tasks such as the NLVR2 benchmark that we mentioned in the opening paragraph and we do not address generative tasks like image captioning here. In our first set of experiments (Section 4), we ask whether it is possible to save compute by freezing parts of the model during pretraining and examining the effect on downstream performance. In our second and final set of experiments (Section 5) we compare the performance of a VL transformer based on a pretrained text encoder versus one based on a pretrained vision transformer. Both sets of experiments will help to establish best training practices for those interested in training VL transformers and hopefully also provide theoretical insight. To perform our experiments, we created a novel VL framework that we call Renaissance that streamlines the ability to evaluate different VL model types (e.g., 1-tower and 2-tower) against a suite of benchmarks.
@@
  • We introduce a software platform Renaissance that offers a range of options for creating, training and testing vision-language transformer encoder models.
@@
2 Related Work
2.1 Pretraining Vision-Language Transformers
The domain of vision-language modeling has seen major advancements in recent years with the adaptation of the transformer (Vaswani et al. 2017a) as VL models. The first examples of VL transformers to appear in the literature were adaptations of the popular BERT NLP model (Devlin et al. 2018). Some examples include VilBERT (Lu et al. 2019), LXMERT (Tan and Bansal 2019a) and VisualBERT (Li et al. 2019a). In the short space of time since these models were introduced a bewildering array of model variations have appeared in the literature. There are huge models designed for zero-shot inference such as Flamingo (Alayrac et al. 2022) and versatile models such as OFA (Wang et al. 2022) that can generate both text and images.
@@ 2.1 Pretraining Vision-Language Transformers
While the literature is now replete with vision-language models, the analysis of their performance and the establishment of best practices has been mostly left open. The aformentioned study Bugliarello et al. (2021) examines pretraining of vision-language model. Bugliarello et al. (2023) is an effort by the same lead author, Emanuele Bugliarello, that provides an analysis of several models on what they term "fine-grained" tasks. Frank et al. (2021) examined the extent to which the vision and language modalities are actually integrated in VL transformers. As valuable as these studies have been however, they have barely scratched the surface of understanding vision-language transformers.
2.2 Vision-Language Modeling Software
@@
Vision language modeling has only recently come to prominence and the available software for it is still in a fairly primitive state. When using NLP models, researchers have a range of available software options that abstract many of the most difficult elements away from users. The most prominent example of this is the Huggingface model hub that specializes in NLP transformers. Though there are a few vision-language models available on Huggingface, there aren’t many and they don’t lend themselves to the modifications that research often demands. In addition to the Huggingface Hub, there have been some efforts toward creating software platforms primarily dedicated to multimodal modeling. LAVIS, introduced in Li et al. (2023) by Salesforce, is one such platform. Though well programmed and relatively straight forward to use, this program offers support very few VL models. Further more, implementing VL tasks is also a fairly involved task. The paucity of available software options led us to create the Renaissance platform for VL modeling that we introduce in the next section.
3 Renaissance: A Versatile Vision-Language Modeling Platform
@@
We now describe the Renaissance program that we use to complete all of the experiments in this study. Because this is its first introduction, we will provide an extensive description of the program and its capabilities. In this section we also take the opportunity to introduce the pretraining tasks, finetuning tasks and the architectural elements required to understand the experimental procedures.
3.1 Capabilities
@@
In this section we describe the capabilities and various options available from the Renaissance platform. The most salient feature of the platform is its ability to easily change the basic architectural features of multi-modal transformers, then train and test them. By simply editing a configuration file, a user can choose a pretrained text encoder or a pretrained vision encoder from the Huggingface hub to insert into the model. In addition to the various architectural options, there are also a number of pretraining and fine-tuning tasks and options available. We will describe these in subsections below.
3.1.1 Model Types
@@
A one-tower encoder model consists of an embedding layer, and a single transformer encoder module followed by a classification layer. Previous examples of one-tower encoders include models such as UNITER Chen et al. (2020) and VisualBERT Li et al. (2019b). In principle, one-tower encoders are very much like NLP encoders such as BERT Devlin et al. (2019) and ELECTRA Clark et al. (2020) with some adaptations for the vision language domain.
One of the key adaptations is that a vision-language model’s embedding layer must accommodate both textual and visual features. For NLP models such as BERT, the embedding layer consists of a single large matrix where each column in the matrix is a vector representing the words in the model’s vocabulary. While one-tower vision-language encoders also have this feature, they have additional components that can process an image into a sequence of vectors. In the current version of the program, the embedding layer will always consist of BERT-style word-piece embeddings for text Devlin et al. (2019) and patch embeddings for image features. Patch embeddings were first introduced as part of the ViT model in the paper Dosovitskiy et al. (2020). Here an image is split in small square patches, the patch is then flattened into a vector and projected to the correct embedding dimension. A visual depiction of patch embeddings from the paper that introduced them can be found in Figure 2. In future versions of the program we hope to include support for using grid features derived from a convolutional neural network to embed images Huang et al. (2020).
The second major component of one-tower encoder models is the transformer encoder stack. The encoder stack for vision-language models is architecturally the same as those found in NLP transformers. Here the only major difference is that encoder’s weights are derived from training on vision language tasks. Renaissance supports the use of the use of most text models on the hub as encoder modules and select variety of vision transformer models on the hub. Specifically, vision models based on the transformer Vaswani et al. (2017b), such as ViT Dosovitskiy et al. (2020), DeiT Touvron et al. (2021), DINO Caron et al. (2021) or BeIT Bao et al. (2021). Convolutional models such as ResNet He et al. (2015), and hybrid models such as ConvNeXT Liu et al. (2022) are not supported. Finally, the classification layer is no different than those found in any other deep learning model. They consist of one or two linear layers that output a score for each possible outcome in the target distribution.
@@
Figure 2: A depiction of patch embeddings from Dosovitskiy et al. (2020).
@@
A two-tower encoder model consists of a text-transformer model, a vision-transformer model and set of cross-modal layers that combine the output of each model into a multimodal feature using cross-attention Lu et al. (2019). In cross-attention layers, the key and value vectors from the visual stream are passed to the multi-head attention mechanism of the textual stream. The key and value vectors from the text stream are also passed to the attention heads in the visual stream resulting in a multi-modal output. Figure 3 shows a simple visual representation of a two-tower model. Because the vision and text modules are separate and the vision and text streams only interact in the cross-modal layers. This is in contrast to one-tower models where visual and textual features interact throughout the model. Another distinction is that for two-tower models the visual and textual features need not be embedded in the same vector space because each encoder module is associated with its own embedding layer. Some previously introduced examples of two-tower transformers are METER Dou et al. (2022) and BridgeTower Xu et al. (2023).
In previously released two-tower encoders, the text encoder modules are architecturally much like BERT and the vision modules much like ViT Dosovitskiy et al. (2020). Renaissance allows users to create new tow-tower models with most vision transformers on the hub as a vision module (convolutional models are not supported), and most text transformers on the hub as a text module. The layers in the cross-modal module are based on implementations from the LXMERT model Tan and Bansal (2019b). Users can choose the dimension and number of cross-modal layers and the number of attention heads per layer. Finally the classification layer is essentially the same as those found in one-tower models.
3.1.2 Training and Configuration Options
Beyond providing flexibility in basic architecture design, the program also provides several options for training and configuring models. The most salient of these features are discussed in the this subsection.
@@ 3.1.2 Training and Configuration Options
Multi-modal models are often initialized with weights from pretrained text or image models. For instance VisualBERT is initialized with the weights from the text model BERT Devlin et al. (2019) and ViLT with weights from the image transformer ViT Dosovitskiy et al. (2020). When doing research it is often useful to initialize model weights randomly and train from scratch. This is often useful for establishing baselines in experiments and as we show in Section 5, can have beneficial to the performance of one-tower models. Users can randomly initialize encoder weights by simply changing settings in a configuration file.
@@
By default, the dimensions of encoder modules is determined by the huggingface hub. However, when model weights are set to be randomly initialized users can manually specify the dimensions of encoder modules. This allows users to easily create completely novel architectures. As an example, consider a one-tower encoder where the encoder is based on ELECTRA-Small. By default, ELECTRA has a hidden size of 256, an embedding size of 128, an intermediate size of 1028 and 12 layers. Any of these numbers can be altered to create encoders of the desired shape and size.
@@
3.1.3 Pretraining Tasks
Currently, our program supports two pretraining tasks, masked language modeling and image-text matching. Models can be pretrained with either of these tasks individually or both in conjunction. Using both tasks in conjunction is a common approach found in the literature. Both tasks are briefly described in the list immediately below. A more thorough description can be found in Fields and Kennington (2023).
  • Masked language modeling (MLM) tasks the model with guessing a masked word based on the image features and the unmasked words. The MLM task was first introduced in Devlin et al. (2019). In the original task, the model’s prediction is based only on the unmasked words in sequence of text. In the multimodal setting, the model’s prediction is based on the unmasked words as well as the associated image.
@@ 3.1.3 Pretraining Tasks
At the time of this writing, Renaissance currently supports four different multimodal datasets for pretraing. These four are Visual Genome Krishna et al. (2017), MSCOCO Lin et al. (2014), Conceptual Captions Sharma et al. (2018) and SBU Captions dataset Ordonez et al. (2011). These can also be used by themselves or in any combination for pretraining models.
3.1.4 Downstream Vision-Language Tasks
In order to test and evaluate models, renaissance currently has five downstream vision-language tasks implemented. They are listed below with a brief description of each task.
@@ 3.1.4 Downstream Vision-Language Tasks
    NLVR2 NLVR2 stands for Natural Language Reasoning for Real and was introduced in Suhr et al. (2018). Here a model will be given two images and must answer a true or false question about them. The addition of second image also makes this quite a challenging task. NLVR2 is very commonly used to benchmark VL models. An example from the dataset can be seen in Figure 4.
@@
    SNLI-VE In the SNLI-VE task is a model is presented with an image text-pair and must determine if the image entails the sentence, contradicts the sentence or is neutral with respect to the sentence. It was introduced in Xie et al. (2019). This task tends to be less challenging than the previous and requires less time to fine-tune and evaluate. Though it appears less commonly in the literature, its quick training time makes it very useful as a model development tool.
@@
    Reference Resolution with RefCOCO In this final task a model is presented with an image that is segmented into several objects and a sentence describing one of these objects. The model must then determine which object the sentence is referring to. The RefCOCO dataset was introduced in Kazemzadeh et al. (2014).
@@
    Multimodal Retrieval with MSCOCO and Flickr30k Multimodal retrieval tasks correspond to activities such as an internet image search. Here a model is given a string of text and must rank a number of images according to how relevant they are to the sentence. The converse process, an image is provided and the model must rank a series of sentences, is also implemented. Our program supports fine-tuning and evaluating both retrieval tasks on the MSCOCO Lin et al. (2014) and Flickr30k datasets.
@@
    Visual Question Answering In this task a model is presented with an image and a question pair and must choose the correct answer from a given set possible choices or generate a free-form answer. The visual question answering task (also called VQA) was introduced in Antol et al. (2015). This task is commonly benchmark in the vision-language field and is also quite challenging. An example from the dataset can be seen in Figure 7.
Figure 4: An example from the NLVR2 dataset. Image from .
Figure 5: An example from the RefCOCO dataset. Image from .
Figure 6: An example from the RefCOCO dataset. Image from .
Figure 7: An example for Visual Question Answering from Suhr et al. (2018).
3.1.5 Unimodal Downstream Tasks
Renaissance also supports downstream evaluation on pure NLP tasks and pure computer vision tasks. For pure NLP, Renaissance supports the GLUE tasks Wang et al. (2018). GLUE is a set of natural language understanding tasks commonly used to benchmark NLP models. The program also supports image classification on the CIFAR10 Krizhevsky et al. (2009). The unimodal capability will be useful in testing if and how multimodal training affects unimodal tasks.
3.2 Design and Implementation
Renaissance is entirely written in the Python programming language. Though popular and user friendly, using and maintaining large-scale python programs can be a difficult process. In order to make this model platform useful as a research tool, we’ve made a number of conscious decisions designed to improve the usability and versatility of the program. These design goals are discussed in the section below.
Modularity
This project grew out of our efforts in designing and training compact VL models. In general, this type of work requires pretraining a large number model architectures with various hyperparameters and finetuning them on suitable evaluation tasks. This type of research demands versatility which indicated a modular design. The program incorporates text and vision models derived from the Huggingface Model Hub as encoder modules in custom vision-language models. Because models from Huggingface are generally written as discrete classes with common methods, this modular approach works quite well. This allows for the user to create models with a wide variety of architectures by simply specifying which models from the hub they would like to use.
Beyond pretrained encoder modules, the other parts of the models are also contained in discrete classes that can be easily substituted. The embedding layer of one-tower models and the cross-modal encoders of two-tower models and all classification modules are written as self-contained classes to allow for easy modification. For instance, to add a new type of classification head, one would simply have to write a new class with compatible methods to the appropriate files.
Ease of Use
Correctly installing dependencies to match particular Python installations and hardware setups can be a major obstacle to using deep learning repositories. In order to make Renaissance as user-friendly as possible we have intentionally tried to reduce to the number of dependencies used in this software platform. Where possible we have used packages that are native to base Python, such as using the PILLOW module for image processing instead of third-party libraries like OpenCV.
Admittedly, the complexity of the underlying processes makes configuring models with this program somewhat difficult. In order facilitate this process, we make use the python module Sacred to track the various model settings and hyperparameters. Sacred assists users to configure models and vary settings to easily reproduce experiments. In addition to this, we provide extensive documentation to make using the program, and importantly extending it for novel purposes, as straightforward a process as possible.
Scalability
Though much of our research is primarily focused on modeling with limited compute resources, the program design is intended to make its use with larger compute setups easy to accommodate. To this end, its models are implemented using PytorchLightning. The PytorchLightning package wraps around Pytorch’s data distributed parallel (DDP) library and abstracts many of the difficult parallel programming aspects away from the user. This makes adding additional nodes or devices much more straightforward than it otherwise would be. Additionally, all of the local data processing functions are handled using the fast and memory efficient PyArrow library.
4 Experiment 1: Freezing Encoder Modules During Pretraining
4.1 Premise

### v2

# Renaissance: A Framework for Investigating the Pretraining of Vision-Language Encoders

@@ Abstract
In the past several years there has been an explosion of available models for vision-language (VL) tasks. Unfortunately, the literature still leaves open a number of questions related to best practices in designing and training such models. Additionally, the limited programming tools available for modeling make conducting VL research more difficult than necessary. In this paper, we seek to answer several questions related to the pretraining of VL encoders through meta-analysis. To conduct these experiments, we introduce a VL evaluation framework called Renaissance. In our first set of experiments, we show that we can save significant compute at little to no cost to downstream performance, by freezing large parts of VL models during pretraining. In our second set of experiments, we examine the effect of basing a VL transformer on a vision model versus a text model. Renaissance offers a great deal of flexibility in creating, training and evaluating transformer encoders for VL modeling. Its source code will be made publicly available upon publication.
Keywords: keyword1, keyword2, keyword3
[…]
Renaissance: A Framework for Investigating the Pretraining of Vision-Language Encoders
Clayton Fields, Casey Kennington
Department of Computer Science
Boise State University
{claytonfields, caseykennington}@boisestate.edu
Abstract content
1.  Introduction
@@
In the span of a few years, dozens of vision-language (VL) transformers have appeared in the literature with a bewildering array of architectures and training methods (see Fields and Kennington (2023b) for a review). VL tasks, such as NLVR2 (Suhr et al., 2018) where the model is tasked with answering questions about images require models to somehow represent and fuse both text and image information. Over the past few years, research in this field has been dominated by the study of generative, decoder-based models. Though generative models offer exciting possibilities and are required for tasks such as image captioning, encoder models are still well suited to classification tasks and are generally smaller and require significantly less compute for training and inference. These characteristics make encoder models a valuable asset for practical applications and for researchers who don’t have access to large-scale compute resources. Unfortunately, as the focus of most research has shifted to very large, decoder-based models, the study of VL encoders has been neglected. This stands in contrast to the natural language procesing (NLP) domain, where studies such as Rogers et al. (2021) have thoroughly investigated the inner workings and best training practices for NLP transformer encoders. To date, there have been only a handful of studies analyzing VL-transformer encoders, such as Bugliarello et al. (2021), and the collected literature still fails to address some very basic questions concerning VL modeling with transformers.
@@
In this paper, we begin to address this gap by providing a systematic analysis geared toward shedding light on some basic aspects of training transformer encoders for VL modeling. As our focus is on the investigation of encoder models, we do not address generative tasks like image captioning here. In our first set of experiments (Section 4), we ask whether it is possible to minimize compute requirements by freezing parts of the model during pretraining and examining the effect on downstream performance. In our second and final set of experiments (Section 5) we compare the performance of a VL transformer based on a pretrained text encoder versus one based on a pretrained vision transformer. Both sets of experiments help establish best training practices for those interested in training VL transformers and hopefully also provide theoretical insight. To perform our experiments, we created a novel VL framework that we call Renaissance that streamlines the ability to evaluate different VL model types (e.g., one-tower and two-tower, explained further in Section 3.1.1) against a suite of benchmarks. This program offers a variety of training options and we hope that it proves a valuable tool for researchers.
@@
  • We introduce a software framework Renaissance that offers a range of options for creating, training and testing VL transformer encoder models.
@@
2.  Related Work
2.1.  Pretraining Vision-Language Transformers
@@
While the literature is now replete with VL models, the analysis of their performance and the establishment of best practices has mostly been applied to large decoder-based, generative models. Some examples include Aghajanyan et al. (2023), McKinzie et al. (2024) and Tong et al. (2024). While the studies offer a great deal of insight into generative models, those interested in using light-weight encoder models to achieve their goals are left with a number of questions unanswered. Some work has been devoted to the analysis of transformer encoders. Bugliarello et al. (2021) examined pretraining of VL model and Bugliarello et al. (2023) provides an analysis of several models on what they term “fine-grained" tasks. Frank et al. (2021) examined the extent to which the vision and language modalities are actually integrated in VL transformers. As valuable as these studies have been, basic questions about VL transformers and their training remain open. These include: which architectures perform best in which situations, best practices for pretraining and what effects pretrained NLP and CV models have on downstream VL performance. This study begins to address some of these.
$2.$  Vision-Language Modeling Framework
@@
When using language models, researchers have a range of available software options that abstract many of the most difficult elements away from users. The most prominent example of this is the Huggingface model hub that specializes transformers. The Huggingface hub has many different model types devoted to NLP and several to computer vision (CV). Unfortunately, there are only a few VL models available and they don’t easily lend themselves to the modifications that research often demands. In addition to the Huggingface Hub, there have been some efforts toward creating software frameworks primarily dedicated to multimodal modeling. LAVIS, introduced in Li et al. (2023) by Salesforce, is one such platform. Though well programmed and relatively straight forward to use, this program supports relatively few VL models. To date, there is still a great deal of space for additional programming options in the field of vision language modeling, particularly for purpose of conducting research. To help fill this space, we created the Renaissance framework for VL modeling that we introduce in the next section.
3.  Renaissance: A Versatile Vision-Language Modeling Framework
@@
In this section, we describe the Renaissance framework that we use to complete all of the experiments in this study. Because this is its first introduction, we provide an extensive description of the framework and its capabilities. We also introduce the pretraining tasks, finetuning tasks and the architectural elements required to understand the experimental procedures.
3.1.  Functionality
@@
In this section we describe the functionality and various options available from the Renaissance framework. The most salient feature of the framework is its ability to easily change the basic architectural features of multi-modal transformers, then train and test them. By simply editing a configuration file, a user can choose a pretrained text encoder or a pretrained vision encoder from the Huggingface hub to insert into the model. In addition to the various architectural options, there are also a number of pretraining and fine-tuning tasks and options available.
3.1.1.  Model Types
@@
A one-tower encoder model consists of an embedding layer, and a single transformer encoder module followed by a classification layer. Previous examples of one-tower encoders include models such as UNITER (Chen et al., 2020) and VisualBERT (Li et al., 2019). In principle, one-tower encoders are very much like NLP encoders such as BERT (Devlin et al., 2019) and ELECTRA (Clark et al., 2020) with some adaptations for the vision language domain.
@@
One of the key adaptations is that a VL model’s embedding layer must accommodate both textual and visual features. For NLP models such as BERT, the embedding layer consists of a single large matrix where each column in the matrix is a vector representing the words in the model’s vocabulary. While one-tower VL encoders also have this feature, they have additional components that can process an image into a sequence of vectors. In the current version of the program, the embedding layer will always consist of BERT-style word-piece embeddings for text (Devlin et al., 2019) and patch embeddings for image features. Patch embeddings were first introduced as part of the ViT model in Dosovitskiy et al. (2020). Here an image is split in small square patches, the patch is then flattened into a vector and projected to the correct embedding dimension. In future versions of the program we hope to include support for using grid features derived from a convolutional neural network to embed images (Huang et al., 2020).
The second major component of one-tower encoder models is the transformer encoder stack. The encoder stack for VL models is architecturally the same as those found in NLP transformers. Here, the only major difference is that encoder’s weights are derived from training on VL tasks. Renaissance supports the use of most text models on the hub as encoder modules and select variety of vision transformer models on the hub. Specifically, vision models based on the transformer (Vaswani et al., 2017), such as ViT (Dosovitskiy et al., 2020), DeiT (Touvron et al., 2021), DINO (Caron et al., 2021) or BeIT (Bao et al., 2021). Convolutional models such as ResNet (He et al., 2015), and hybrid models such as ConvNeXT (Liu et al., 2022) are not supported for one-tower modeling. Finally, the classification layer is no different than those found in any other deep learning model. They consist of one or two linear layers that output a score for each possible outcome in the target distribution.
@@
A two-tower encoder model consists of a text-transformer model, a vision model and set of cross-modal layers that combine the output of each model into a multimodal feature using cross-attention (Lu et al., 2019). In cross-attention layers, the key and value vectors from the visual stream are passed to the multi-head attention mechanism of the textual stream. The key and value vectors from the text stream are also passed to the attention heads in the visual stream resulting in a multi-modal output. Figure 2 shows a simple visual representation of a two-tower model. Because the vision and text modules are separate and the vision and text streams only interact in the cross-modal layers. This is in contrast to one-tower models where visual and textual features interact throughout the model. Another distinction is that for two-tower models the visual and textual features need not be embedded in the same vector space because each encoder module is associated with its own embedding layer. Some previously introduced examples of two-tower transformers are METER (Dou et al., 2022) and BridgeTower (Xu et al., 2023).
In previously released two-tower encoders, the text encoder modules are architecturally much like BERT and the vision modules much like ViT (Dosovitskiy et al., 2020). Renaissance allows users to create new two-tower models with most of the vision modules on the hub as a vision module. A novel feature of this framework is that users can also incorporate convolutional models such as ResNet (He et al., 2015) and hybrid models such as ConvNext (Liu et al., 2022). Most text transformers on the hub can be used as text modules in two-tower models. The layers in the cross-modal module are based on the implementation from the LXMERT model (Tan and Bansal, 2019). Users can choose the dimension and number of cross-modal layers and the number of attention heads per layer. Finally, the classification layer is essentially the same as those found in one-tower models.
3.1.2.  Training and Configuration Options
Beyond providing flexibility in basic architecture design, the program also provides several options for training and configuring models. The most notable of these features are discussed in this subsection.
@@ 3.1.2.  Training and Configuration Options
Multi-modal models are often initialized with weights from pretrained text or image models. For instance VisualBERT is initialized with the weights from the text model BERT (Devlin et al., 2019) and ViLT with weights from the image transformer ViT (Dosovitskiy et al., 2020). When doing research it is often useful to initialize model weights randomly and train from scratch. This is useful for establishing baselines in experiments and as we show in Section 5, can be beneficial to the performance of one-tower models. Users can randomly initialize encoder weights by simply changing settings in a configuration file.
@@
By default, the dimensions of encoder modules is determined by the huggingface hub. However, when model weights are set to be randomly initialized users can manually specify the dimensions of encoder modules. This allows users to easily create completely novel architectures. As an example, consider a one-tower encoder where the encoder is based on ELECTRA-Small. By default, ELECTRA has a hidden size of 256, an embedding size of 128, an intermediate size of 1028 and 12 layers. Any of these hyperparameters can be altered to create encoders of the desired shape and size.
@@
3.1.3.  Pretraining Tasks
Currently, our program supports two pretraining tasks, masked language modeling and image-text matching. Models can be pretrained with either of these tasks individually or both in conjunction. Using both tasks in conjunction is a common approach found in the literature. Both tasks are briefly described in the list immediately below. A more thorough description can be found in Fields and Kennington (2023b).
  • Masked language modeling (MLM) tasks the model with guessing a masked word based on the image features and the unmasked words. The MLM task was first introduced in Devlin et al. (2019). In the original task, the model’s prediction is based only on the unmasked words in sequence of text. In the multimodal setting, the model’s prediction is based on the unmasked words as well as the associated image features.
@@ 3.1.3.  Pretraining Tasks
At the time of this writing, Renaissance currently supports four different multimodal datasets for pretraining. These four are Visual Genome (Krishna et al., 2017), MSCOCO (Lin et al., 2014), Conceptual Captions (Sharma et al., 2018) and SBU Captions dataset (Ordonez et al., 2011). They can also be used by themselves or in any combination for pretraining models. These four datasets are commonly used in the pretraining of VL models.
3.1.4.  Downstream Vision-Language Tasks
In order to test and evaluate models, renaissance currently has five downstream VL tasks implemented. They are listed below with a brief description of each task.
@@ 3.1.4.  Downstream Vision-Language Tasks
    NLVR2 NLVR2 stands for Natural Language Reasoning for Real and was introduced in Suhr et al. (2018). Here a model will be given two images and must answer a true or false question about them. The addition of second image also makes this quite a challenging task. NLVR2 is very commonly used to benchmark VL models.
@@
    SNLI-VE In the SNLI-VE task is a model is presented with an image text-pair and must determine if the image entails the sentence, contradicts the sentence or is neutral with respect to the sentence. It was introduced in (Xie et al., 2019). This task tends to be less challenging than the previous and requires less time to fine-tune and evaluate. Though it appears less commonly in the literature, its quick training time makes it very useful as a model development tool.
@@
    Reference Resolution with RefCOCO In this final task a model is presented with an image that is segmented into several objects and a sentence describing one of these objects. The model must then determine which object the sentence is referring to. The RefCOCO dataset was introduced in (Kazemzadeh et al., 2014).
@@
    Multimodal Retrieval with MSCOCO and Flickr30k Multimodal retrieval tasks correspond to activities such as an internet image search. Here a model is given a string of text and must rank a number of images according to how relevant they are to the sentence. The converse process, an image is provided and the model must rank a series of sentences, is also implemented. Our program supports fine-tuning and evaluating both retrieval tasks on the MSCOCO (Lin et al., 2014) and Flickr30k datasets.
@@
    Visual Question Answering In this task a model is presented with an image and a question pair and must choose the correct answer from a given set possible choices or generate a free-form answer. The visual question answering task (also called VQA) was introduced in Antol et al. (2015). This task is a commonly used benchmark in the VL field and is also quite challenging.
Figure 3: An example from the NLVR2 dataset. Image from Suhr et al. (2018).
4.  Experiment 1: Freezing Encoder Modules During Pretraining
4.1.  Purpose